# FedSpaLLM: Federated Pruning of Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) achieve state-of-the-art performance but are challenging to deploy due to their high computational and storage demands. Pruning can reduce model size, yet existing methods assume public access to calibration data, which is impractical for privacy-sensitive applications. To address the challenge of pruning LLMs in privacy-preserving settings, we propose FedSpaLLM, the first federated learning framework designed specifically for pruning LLMs. FedSpaLLM enables clients to locally prune their models based on private data while accounting for system heterogeneity and maintaining communication efficiency. Our framework introduces several key innovations: (1) a novel $\ell_{0}$-norm aggregation function that ensures only non-zero weights are averaged across clients, preserving important model parameters; (2) an adaptive mask expansion technique that meets global sparsity targets while accommodating client-specific pruning decisions; and (3) a layer sampling strategy that reduces communication overhead and personalizes the pruning process based on client resources. Extensive experiments show that FedSpaLLM improves pruning performance in diverse federated settings. Source code will be released upon publication.
@@ 2 Related Work
Federated Learning with LLMs. FL on LLMs has primarily focused on LLM fine-tuning and has gained attention for enabling private and efficient model updates. FedPrompt (Zhao et al., 2023) introduces prompt-tuning in FL, reducing communication costs by updating only soft prompts. FedPipe (Fang et al., 2024) leverages low-rank adapters for parameter-efficient fine-tuning, improving training speed and accuracy. FedNLP (Lin et al., 2022) provides a benchmarking framework for evaluating FL methods on NLP tasks. FedAdapter (Cai et al., 2023) uses adapters to accelerate model convergence in federated settings. FeDeRA (Yan et al., 2024) employs singular value decomposition to further improve LoRA-based fine-tuning efficiency. C2A (Kim et al., 2023) introduces a hypernetwork-based framework for generating client-specific adapters to handle client heterogeneity. FedBPT (Sun et al., 2024) enables efficient prompt-tuning with a gradient-free approach, reducing memory and communication costs. PrE-Text (Hou et al., 2024) generates differentially private synthetic data to enable central training, reducing on-device computation.
Federated Learning with Pruning on DNNs. Model pruning in FL improves efficiency by reducing communication and computation costs. FedP3 (Yi et al., 2024) and HeteroFL (Diao et al., 2021) address client model heterogeneity, enabling smaller, personalized models. FedTiny (Huang et al., 2023) and PruneFL (Jiang et al., 2022) implement progressive pruning to fit models within resource constraints. FedPrune (Munir et al., 2021) improves performance by pruning global models based on client capabilities, while Complement Sparsification (Jiang and Borcea, 2023) reduces communication overhead using sparsity techniques. However, all the work above only applies to smaller DNNs, and cannot trivially scale to massive LLMs.
@@ 3.2 Background of Federated Learning
FL is a distributed machine learning paradigm designed to enable collaborative training of a single global model without exchanging private data between clients. In this setup, multiple clients, each with their own local dataset, train their local models independently. The central server coordinates the process by sending the global model to a selected group of clients. These clients then perform local optimization using their respective datasets. After local training, the updated models are sent back to the server, where they are aggregated to update the global model.
@@
There exist various methods for aggregating the client updates, with FedAvg being the most widely adopted due to its simplicity and effectiveness in a wide range of FL applications. This method aggregates the client updates weighted by their respective dataset sizes, providing a robust way to train models in data-sensitive environments.
@@ 4.1 Problem Formulation
Our formulation is general and supports model heterogeneity, where each client may have its own model parameters, $W_{i}$, as opposed to the vanilla FL setting with sharing a common global model. This flexibility is crucial for accommodating diverse client environments, making our method applicable to a wide range of real-world federated pruning scenarios.
@@ 4.2 Aggregation Function
$C_{j}=\sum\nolimits_{i=1}^{N}\mathbb{I}(M_{i}^{(j)}=0),$ (6)
$k=\text{ceil}(s\cdot d)-\sum\nolimits_{j=1}^{d}\mathbb{I}(C_{j}=N),$ (7)
@@ 5 Experiments
We implement our FedSpaLLM in PyTorch Paszke et al. (2019) and use the HuggingFace Transformers library Wolf et al. (2019) for handling models and datasets. All pruning experiments are conducted on NVIDIA A100 GPUs. For each client, we utilize SparseGPT Frantar and Alistarh (2023) to perform pruning. For the calibration data, we follow Frantar and Alistarh (2023) and use 2048-token segments, randomly chosen from the first shard of the C4 Raffel et al. (2020) dataset. This represents generic text data crawled from the internet and ensures that our experiments remain zero-shot since no task-specific data is seen during pruning.
In particular, we perform experiments on the OPT model family Zhang et al. (2022) and LlaMA-2 model family Touvron et al. (2023) with 4, 8, and 16 clients. We consider OPT-125m, OPT-1.3b, and LlaMA-2 7b with unstructured sparsity of 50% to 80%, and all clients share a copy of the same global model. Each client is assumed to have 32 calibration samples to perform its own pruning and we perform one round of communication. These models of different sizes together with different numbers of clients can provide a broader picture of the performances of FedSpaLLM. In terms of metrics, we mainly focus on perplexity, which is known to be a challenging and stable metric that is well-suited for evaluating the accuracy of compression methods Yao et al. (2022); Dettmers and Zettlemoyer (2023), and thus measuring the performances of the compressed models. We consider the test sets of raw-WikiText2 Merity et al. (2016) (WT-2) and PTB Marcus et al. (1994) as well as a subset of the C4 validation data, which are all popular benchmarks in LLM compression literature Yao et al. (2022); Park et al. (2022); Frantar and Alistarh (2023); Sun et al. (2023).
@@ 5.1 Results and analyses
We present the perplexity results in Tables 1 to 3. In the tables, we report the average perplexity of the client models, denoted by “standalone", and the global model, denoted by “FedSpaLLM". From the results, we the global model from the proposed FedSpaLLM achieves better perplexity across the considered models and levels of sparsity. In general, as the target sparsity increases, we see more significant improvements in the perplexity of the global model over the client models. This is expected because as the sparsity increases, more information is required to accurately prune the model weights to maintain its performance. In essence, all the client models are pruned with the private calibration samples of each client while the global model benefits from the information in the communicated weights from the clients, which can be thought of as having a much larger calibration sample set. Note that in FL setting, such a centralized calibration sample set is not allowed as the client’s calibration samples are private. In addition, we can see the improvements in the perplexity of the global model over the client models are more significant for the LlaMA-2 model family and the PTB dataset. We observe that, in the case of LlaMA-2 7b, the client models generally struggle with the PTB dataset from sparsity 60% and beyond, regardless of the number of clients. In many of those cases, the global model achieves up to 5 times better perplexity. This demonstrates the effectiveness of our proposed FedSpaLLM framework. In addition, we observe there is no obvious trend in perplexity when the number of clients change. Perplexities of the global models are very comparable for different number of clients in FL, when sparsity is small.

### v2

@@ Abstract
Large Language Models (LLMs) achieve state-of-the-art performance but are challenging to deploy due to their high computational and storage demands. Pruning can reduce model size, yet existing methods assume public access to calibration data, which is impractical for privacy-sensitive applications. To address the challenge of pruning LLMs in privacy-preserving settings, we propose FedSpaLLM, the first federated learning framework designed specifically for pruning LLMs. FedSpaLLM enables clients to locally prune their models based on private data while accounting for system heterogeneity and maintaining communication efficiency. Our framework introduces several key innovations: (1) a novel $\ell_{0}$-norm aggregation function that ensures only non-zero weights are averaged across clients, preserving important model parameters; (2) an adaptive mask expansion technique that meets global sparsity targets while accommodating client-specific pruning decisions; and (3) a layer sampling strategy that reduces communication overhead and personalizes the pruning process based on client resources. Extensive experiments show that FedSpaLLM improves pruning performance in diverse federated settings. The source code can be found at https://github.com/BaiTheBest/FedSpaLLM.
@@ 2 Related Work
Federated Learning with LLMs. FL on LLMs has primarily focused on LLM fine-tuning and has gained attention for enabling private and efficient model updates. FedPrompt (Zhao et al., 2023) introduces prompt-tuning in FL, reducing communication costs by updating only soft prompts. Fang et al. (2024) and Li et al. (2024b) leverage low-rank adapters for parameter-efficient fine-tuning, improving training speed and accuracy. FedNLP (Lin et al., 2022) provides a benchmarking framework for evaluating FL methods on NLP tasks. FedAdapter (Cai et al., 2023) uses adapters to accelerate model convergence in federated settings. FeDeRA (Yan et al., 2024) employs singular value decomposition to further improve LoRA-based fine-tuning efficiency. C2A (Kim et al., 2023) introduces a hypernetwork-based framework for generating client-specific adapters to handle client heterogeneity. FedBPT (Sun et al., 2024) enables efficient prompt-tuning with a gradient-free approach, reducing memory and communication costs. PrE-Text (Hou et al., 2024) generates differentially private synthetic data to enable central training, reducing on-device computation.
Federated Pruning on DNNs. Model pruning in FL improves efficiency by reducing communication and computation costs. FedP3 (Yi et al., 2024) and HeteroFL (Diao et al., 2021) address client model heterogeneity, enabling smaller, personalized models. FedTiny (Huang et al., 2023) and PruneFL (Jiang et al., 2022) implement progressive pruning to fit models within resource constraints. FedPrune (Munir et al., 2021) improves performance by pruning global models based on client capabilities, while Complement Sparsification (Jiang and Borcea, 2023) reduces communication overhead using sparsity techniques. However, all the work above only applies to smaller DNNs, and cannot trivially scale to massive LLMs.
Federated Distillation. In addition to model pruning, knowledge distillation (KD) has been explored as a resource-efficient approach to training NNs in edge environments Li et al. (2024a). In this paradigm, a smaller proxy model is used to facilitate learning, either by mimicking the behavior of a larger model or by aggregating knowledge from multiple clients. For instance, FedBiOT Wu et al. (2024) introduces a method for locally fine-tuning LLMs without requiring full-model transmission, allowing clients to distill and learn from a proxy model. Similarly, DaFKD Wang et al. (2023), a domain-aware FL knowledge distillation framework, incorporates domain-specific adaptations to improve the generalization of models across heterogeneous clients.
Despite the merits of federated distillation, our proposed approach offers several advantages. Proxy models and FL distillation methods require training additional models, increasing computational overhead. Additionally, they assume an accessible public dataset for generating the soft labels, which usually does not hold in practice. Our framework directly prunes the global LLM without auxiliary models and any additional public data. It achieves target sparsity across heterogeneous clients while maintaining privacy through local pruning and aggregation.
@@ 3.2 Background of Federated Learning
FL is a distributed machine learning paradigm designed to enable collaborative training of a single global model without exchanging private data between clients Gu et al. (2023). In this setup, multiple clients, each with their own local dataset, train their local models independently. The central server coordinates the process by sending the global model to a selected group of clients. These clients then perform local optimization using their respective datasets. After local training, the updated models are sent back to the server, where they are aggregated to update the global model.
@@
There exist various methods for aggregating client updates, with FedAvg McMahan et al. (2017) being the most widely adopted due to its simplicity and effectiveness in a wide range of FL applications. This method aggregates the client updates weighted by their respective dataset sizes, providing a robust way to train models in data-sensitive environments.
@@ 4.1 Problem Formulation
Our formulation is general and supports model heterogeneity, where each client may have its own model parameters, $W_{i}$, as opposed to the vanilla FL setting sharing a common global model. This flexibility is crucial for accommodating diverse client environments, making our method applicable to a wide range of real-world federated pruning scenarios.
@@ 4.2 Aggregation Function
$C_{j}=\sum\nolimits_{i=1}^{N}\mathbb{I}(M_{i}^{(j)}=0),$ (6)
$k=\text{ceil}(s\cdot d)-\sum\nolimits_{j=1}^{d}\mathbb{I}(C_{j}=N),$ (7)
@@ 4.3 Personalization with Layer Sampling
Theorem 1 ensures that the global model obtained by layer sampling converges to the same result as if every client had pruned all layers. The key here is the randomness of the layer selection process: over multiple communication rounds, the expected contribution of each layer is preserved, resulting in an unbiased estimate of the fully pruned model. This guarantees that our sampling strategy does not introduce systematic errors and that, over time, the sampled model will mirror the performance of the fully pruned model.
@@ 5 Experiments
We implement our FedSpaLLM in PyTorch Paszke et al. (2019) and use the HuggingFace Transformers library Wolf et al. (2019) for handling models and datasets. All pruning experiments are conducted on NVIDIA A100 GPUs. For each client, we utilize SparseGPT Frantar and Alistarh (2023) to perform pruning. For the calibration data, we follow Frantar and Alistarh (2023) and use 2048-token segments, randomly chosen from the first shard of the C4 Raffel et al. (2020) dataset. This represents generic text data crawled from the internet and ensures that our experiments remain zero-shot since no task-specific data is seen during pruning. In addition, we consider a random pruning baseline in which the model is randomly pruned to the target sparsity.
In particular, we perform experiments on the OPT model family Zhang et al. (2022) and LlaMA-2 model family Touvron et al. (2023) with 4, 8, and 16 clients. We consider OPT-125m, OPT-1.3b, and LlaMA-2 7b with unstructured sparsity of 50% to 80%. In each communication round, each of the clients receives a copy of the same global model from the server and each client is assumed to utilize 32 calibration samples to perform its own pruning. By evaluating models of varying sizes alongside different number of clients, we can gain a more comprehensive understanding of FedSpaLLM’s performances in its scalability and robustness. In terms of metrics, we mainly focus on perplexity, which is known to be a challenging and stable metric that is well-suited for evaluating the accuracy of compression methods Yao et al. (2022); Dettmers and Zettlemoyer (2023), and thus measuring the performances of the compressed models. We consider the test sets of raw-WikiText2 Merity et al. (2016) (WT-2) and PTB Marcus et al. (1994) as well as a subset of the C4 validation data, which are all popular benchmarks in LLM compression literature Yao et al. (2022); Park et al. (2022); Frantar and Alistarh (2023); Sun et al. (2023).
@@ 5.1 Results and analyses
We present the perplexity results in Tables 1 to 3. In the tables, we report the random pruning baseline, denoted by “Random", the average perplexity of the client models, denoted by “standalone", and the global model, denoted by “FedSpaLLM". From the results, we see that random pruning results in perplexity that are orders of magnitude higher than both standalone and FedSpaLLM. As the model size increases, the performances of random pruning becomes even worse. This suggests that random pruning may have pruned weights that are crucial for maintaining model quality. Comparing standalone and FedSpaLLM, we see that across the models, datasets, and sparsity levels, FedSpaLLM consistently outperforms standalone in achieving lower perplexity. In general, as the target sparsity increases, we see more noticeable improvements in the perplexity of the global model over the client models. This is expected because as the sparsity increases, more information is required to accurately prune the model weights to maintain the model performances. In essence, each of the client models is pruned with the private calibration samples of each client while the global model benefits from the collaborative information from the communicated weights from the clients. As a result, the global model is effectively utilizing a signicantly larger calibration sample set, even though such a centralized calibration sample set is prohibited in FL setting as the client’s calibration samples are private. Notably, FedSpaLLM achieves substantially lower perplexity compared to standalone in higher sparsity levels, highlighting the benefits of FedSpaLLM where the clients collaboratively contribute to the global model with much better performances while the privacy of their own data is well maintained. Furthermore, we can see the improvements in the perplexity of the global model over the client models are particularly significant for the LlaMA-2 model family and the PTB dataset. We observe that, in the case of LlaMA-2 7b, the client models generally struggle with the PTB dataset from sparsity 60% and beyond, regardless of the number of clients. In many of those cases, the global model achieves up to 5 times better perplexity values. This demonstrates the effectiveness of FL in the pruning tasks. In addition, we do not observe there is noticeable trend in perplexity values with varying number of clients and the perplexity values of the global models are comparable regardless of the number of clients participating in FL, when the sparsity level is small.
@@ 7 Limitations
8 Acknowledgement
This work was supported by the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research, under Contract DE-AC02-06CH11357. This research used resources of the Argonne Leadership Computing Facility at Argonne National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under contract DE-AC02-06CH11357.